OpenStudy (anonymous):

The manager of a basketball arena randomly selected 150 fans and asked if they were pleased with the new selection of food offered in the arena. In this sample, 120 fans said that they were pleased. How many of the 8,000 fans at the game could be expected to be pleased? A. 6400 B. 7850 C. 7880 D. 7970

OpenStudy (kropot72):

The fraction of the fans in the sample who were pleased = 120/150 Therefore the expected number of the 8000 fans at the game who were pleased is given by: \[\large \frac{120}{150}\times\frac{8000}{1}=\frac{120\times8000}{150}\]

OpenStudy (anonymous):

960000

OpenStudy (anonymous):

divide that number by 150 then you get A your welcome futer people.

OpenStudy (kropot72):

Yes, A is the correct choice.

OpenStudy (anonymous):

thanks bunches

OpenStudy (kropot72):

You're welcome :)